# Exploring Lattice Thermal Conductivity Models via Interpretable Deep Learning to Accelerate Materials Discovery

### v8

@@ Abstract
Lattice thermal conductivity, being integral to thermal transport properties, is indispensable to advancements in areas such as thermoelectric materials and thermal management. Traditional methods, such as Density Functional Theory and Molecular Dynamics, require significant computational resources, posing challenges to the high-throughput prediction of lattice thermal conductivity. Although AI-driven material science has achieved fruitful progress, the trade-off between accuracy and interpretability in machine learning continues to hinder further advancements. This study utilizes interpretable deep learning techniques to construct a rapid prediction framework that enables both qualitative assessments and quantitative predictions, accurately forecasting the thermal transport properties of three novel materials. Furthermore, interpretable deep learning offers analytically grounded physical models while integrating with sensitivity analysis to uncover deeper theoretical insights.
@@ I Introduction
Kolmogorov-Arnold Networks (KANs) [30] , a novel neural network architecture distinct from Multi-Layer Perceptrons (MLPs), show significant promise in solving partial differential equations (PDEs), time series prediction, and high-dimensional data tasks [31, 30, 32]. In some cases, smaller KAN models outperform MLPs [31, 30]. KANs have also demonstrated success in engineering applications, such as predicting chiller energy consumption [33] and nuclear binding energy [34], highlighting both feasibility and broad potential. Notably, their capabilities in symbolic regression offer interpretability advantages that absent in MLPs [31, 30]. As previously mentioned, existing work has successfully used black-box models to accurately predict $\kappa_{\rm L}$. However, achieving accurate $\kappa_{\rm L}$ predictions while simultaneously visualizing the model’s decision-making process, akin to white-box model, remains a significant challenge. The emergence of KANs provides an opportunity to address this issue.
Figure 1 illustrates the framework of our work, which aims to model and predict $\kappa_{\rm L}$ using the white-box deep learning model KAN and compare its performance with that of conventional black-box and white-box ML models. In addition to accuracy, we perform sensitivity analysis to examine interpretability differences across various white-box models and to assess the contributions of individual features to $\kappa_{\rm L}$. This approach further elucidates the relationships between $\kappa_{\rm L}$ and key features, as well as feature-feature correlations. Additionally, Crystal Graph Convolutional Neural Network (CGCNN) [35] is employed to predict critical physical features required for $\kappa_{\rm L}$ modeling with high accuracy, forming a two-stage framework that significantly reduces the time required for screening potential materials. This framework enables rapid qualitative assessments of thermal insulators and conductors. Finally, by integrating DFT and MD, our pretrained KAN model achieves outstanding accuracy in predicting the properties of candidates in an unlabeled extrapolation dataset.
@@ II.1 Establishment and evaluation of $\kappa_{\rm L}$ models
Figure 2: Model performance on primary dataset. (a) The blue circular markers represent samples from the training set, while the red star scatters are samples from the test set. (b) KAN model architecture based on the primary dataset.
@@
  • Molecular features are often non-smooth, whereas deep models tend to favor smooth solutions [48, 49];
@@
$\log{(\kappa)_{{\rm{SISSO}}}}=$ $-2.8+0.3\log\frac{{{\Theta^{2}}}}{{\left|{{C_{\rm{p}}}-{C_{\rm{v}}}}\right|}}$
@@
the linear coefficients in the model are obtained through least-square regression. For the specific physical meanings of the features, please refer to Table 2. Our dataset consists of a total of 10 features, but the SISSO model utilizes only seven of these. Some features, such as the space group number, which are typically not considered to be related to $\kappa_{\rm L}$. SISSO effectively identifies and excludes such extraneous variables. However, it overlooks physical quantities that are theoretically linked to $\kappa_{\rm L}$, such as the thermal expansion coefficient [50, 51]. Our feature sensitivity analysis indicates that the thermal expansion coefficient is regarded as an important parameter for modeling $\kappa_{\rm L}$.
@@
Feature Abbreviation Unit Range
@@ II.2 Deepening physical interpretability through sensitivity analysis
where $Q$ represents the total enthalpy of formation of the compound, denotes Kelvin temperature (here $T=300{\rm K}$), and the equation above can be rewritten as $\frac{F}{{{S_{\rm{v}}}}}=\frac{Q}{{{S_{\rm{v}}}}}-T$. Due to the presence of the entropy-enthalpy compensation (EEC) [57] effect, a linear relationship exists between entropy and enthalpy, which is further reflected in the linear correlation between $S_{\rm v}$ and $F$. Based on the work of G. D. Garbulsky and G. Ceder [58], within the harmonic approximation, when the temperature is greater than the system’s characteristic Debye temperature, the lattice Hamiltonian can be written as:
@@
$\log{\left(\kappa\right)_{{\rm{KAN}}}}=$ $-1.05\sin d_{1}+0.61\cosh d_{2}-1.44$
@@
$\log{\left(\kappa\right)_{{\rm{SISSO}}}}=$ $1.53+0.0079\frac{{F\log\left({{C_{\rm{v}}}}\right)}}{{\sqrt[3]{{{C_{\rm{v}}}}}}}$
@@ II.3 Screening potential thermal insulators/conductors
  • Excluding transition metal elements;
  • Number of atoms (nsites) $<$ 5;
  • Band gap $\in\left({0,1.5{\rm{eV}}}\right]$.
Figure 6: $k$-means clustering and t-SNE 2D visualization. (a) Labeled training samples from AFLOW are clustered into seven groups, each represented by a different color. (b) The same scatter plot as in (a), but the samples are colored according to their $\log(\kappa_{\rm L})$, with black pentagrams representing unlabeled candidates from the MP database. (c) and (d) are zoomed-in views showing the thermal insulator/thermal conductor candidates in the t-SNE space.
In Figure 6(a), all training samples are distinctly clustered into seven clusters, $C_{1}$ to $C_{4}$, with well-defined boundaries between clusters. Mapping the values of $\log(\kappa_{\rm L})$ onto the classification results in Figure 6(a), as shown in Figure 6(b), we observe that as t-SNE component 1 increases, $\log(\kappa_{\rm L})$ exhibits an upward trend. This allows us to identify $C_{1}$ as the cluster corresponding to thermal insulators and $C_{4}$ as the cluster representing thermal conductors. More importantly, the clustering results indicate that $\log(\kappa_{\rm L})$ can be accurately and qualitatively assessed using only the three features $F$, $K_{\rm s}$, and $\alpha$. Based on the labeled dataset, we obtained a CGCNN pre-trained model for the initial features and applied it to predict $F$, $K_{\rm s}$, and $\alpha$ for unlabeled samples from the MP database. These predicted features were then clustered in the same manner as the labeled samples. In Figure 6(b), we selected the 10 most promising thermal insulator/conductor candidates from clusters $C_{1}$ and $C_{4}$ (see Table 3). The zoomed-in views are shown in Figures 6(c) and (d), where these candidates are found in the same clusters as the labeled samples with the lowest/highest $\log(\kappa_{\rm L})$, respectively. Therefore, we focus on these 10 candidates for further predictions and validation. Existing literature [75, 76, 77, 78, 79, 80] on these candidates primarily focuses on 2D materials or electronic thermal conductivity. Our research targets the $\kappa_{\rm L}$ of 3D materials, aiming to fill this gap in the field.
Figure 7: Validation of the KAN predictions. For (a) Be2C, (b) CN2 and (c) CrB4, a relaxation time of 5 ns is sufficient for the thermal conductivity to converge. The deep red dashed line represents the thermal conductivity predicted by the KAN model, while the light red band indicates the prediction range with a ±5% error margin. The $\kappa_{\rm L}$ calculated by MD is represented by the blue line, illustrating its convergence behavior with respect to relaxation time. (d) shows a histogram of the thermal conductivity distribution in the primary dataset.
According to our MD calculations, the actual $\kappa_{\rm L}$ values of the candidate materials fall within distinct ranges ($\sim$30, $\sim$80, $\sim$150 W/m/K). As shown in Figure 7(d), samples with $\kappa_{\rm L}>100$ W/m/K account for only 0.47% of the entire training dataset. In such cases of data imbalance, model predictions for the target region with scarce samples typically exhibit significant uncertainty [81, 82]. Nevertheless, the $\kappa_{\rm L}$ values predicted by KAN remain in excellent agreement with the MD validation results. This demonstrates that the pretrained KAN model achieves not only reliable accuracy but also satisfactory extrapolation capability. The methods of $k$-means clustering and t-SNE visualization are also sufficient for qualitatively assessing the $\kappa_{\rm L}$ of materials, enabling high-throughput screening in combination with the pretrained model.
Table 3: The predicted results of $\kappa_{\rm L}$ based on KAN
@@
II.4 Tracing high/low $\kappa_{\rm L}$ from a theoretical perspective

### v9

# Accelerating the Exploration of Thermal Materials via a Synergistic Strategy of DFT and Interpretable Deep Learning

@@ Abstract
Lattice thermal conductivity (LTC) is a critical parameter for thermal transport properties, playing a pivotal role in advancing thermoelectric materials and thermal management technologies. Traditional computational methods, such as Density Functional Theory (DFT) and Molecular Dynamics (MD), are resource-intensive, limiting their applicability for high-throughput LTC prediction. While AI-driven approaches have made significant strides in material science, the trade-off between accuracy and interpretability remains a major bottleneck. In this study, we introduce an interpretable deep learning framework that enables rapid and accurate LTC prediction, effectively bridging the gap between interpretability and precision. Leveraging this framework, we identify and validate four promising thermal conductors/insulators using DFT and MD. Moreover, by combining sensitivity analysis with DFT calculations, we uncover novel insights into phonon thermal transport mechanisms, providing a deeper understanding of the underlying physics. This work not only accelerates the discovery of thermal materials but also sets a new benchmark for interpretable AI in material science.
@@ I Introduction
Kolmogorov-Arnold Networks (KANs) [30] , a novel neural network architecture distinct from Multi-Layer Perceptrons (MLPs), show significant promise in solving ordinary differential equations [31], time series classification [32], and fluid modeling [33]. In some cases, smaller KAN models outperform MLPs [RN165, 30]. Notably, their capabilities in symbolic regression offer interpretability advantages that absent in MLPs [RN165, 30]. As previously mentioned, existing work has successfully used black-box models to accurately predict $\kappa_{\rm L}$. However, achieving accurate $\kappa_{\rm L}$ predictions while simultaneously visualizing the model’s decision-making process, akin to white-box model, remains a significant challenge. The emergence of KANs provides an opportunity to address this issue.
Figure 1 illustrates the framework of our work, which aims to model and predict $\kappa_{\rm L}$ using the white-box deep learning model KAN and compare its performance with that of conventional black-box and white-box ML models. In addition to accuracy, we perform sensitivity analysis to examine interpretability differences across various white-box models and to assess the contributions of individual features to $\kappa_{\rm L}$. This approach further elucidates the relationships between $\kappa_{\rm L}$ and key features, as well as feature-feature correlations. Additionally, Crystal Graph Convolutional Neural Network (CGCNN) [34] is employed to predict critical physical features required for $\kappa_{\rm L}$ modeling with high accuracy, forming a two-stage framework that significantly reduces the time required for screening potential materials. This framework enables rapid qualitative assessments of thermal insulators and conductors. Furthermore, validated through DFT and MD calculations, we successfully identified three thermal conductors and one thermal insulator using this framework. Finally, taking two of these materials as case studies, we leverage DFT to provide in-depth insights into their respective phonon thermal transport mechanisms.
@@ II.1 Establishment and evaluation of $\kappa_{\rm L}$ models
Figure 2: Model performance on primary dataset. (a) The blue circular markers represent samples from the training set, while the red square scatters are samples from the test set. (b) KAN model architecture based on the primary dataset.
@@
  • Chemical features are often non-smooth, whereas deep models tend to favor smooth solutions [47, 48];
@@
$\log{(\kappa_{\rm L})_{{\rm{SISSO}}}}=$ $-2.8+0.3\log\frac{{{\Theta^{2}}}}{{\left|{{C_{\rm{p}}}-{C_{\rm{v}}}}\right|}}$
@@
the linear coefficients in the model are obtained through least-square regression. For the specific physical meanings of the features, please refer to Table 2. Our dataset consists of a total of 10 features, but the SISSO model utilizes only seven of these. Some features, such as the space group number, which are typically not considered to be related to $\kappa_{\rm L}$. SISSO effectively identifies and excludes such extraneous variables. However, it overlooks physical quantities that are theoretically linked to $\kappa_{\rm L}$, such as the thermal expansion coefficient [49, 50], while feature sensitivity analysis indicates that the thermal expansion coefficient is regarded as an important parameter for modeling $\kappa_{\rm L}$ (see Section II.2).
@@
Definition Notation Unit Range
@@ II.2 Deepening physical interpretability through sensitivity analysis
where $Q$ represents the total enthalpy of formation of the compound, denotes Kelvin temperature (here $T=300{\rm K}$), and the equation above can be rewritten as $\frac{F}{{{S_{\rm{v}}}}}=\frac{Q}{{{S_{\rm{v}}}}}-T$. Due to the presence of the entropy-enthalpy compensation [56] effect, a linear relationship exists between entropy and enthalpy, which is further reflected in the linear correlation between $S_{\rm v}$ and $F$. Based on the work of G. D. Garbulsky and G. Ceder [57], within the harmonic approximation, when the temperature is greater than the system’s characteristic Debye temperature, the lattice Hamiltonian can be written as:
@@
$\log{\left(\kappa_{\rm L}\right)_{{\rm{KAN}}}}=$ $-1.05\sin d_{1}+0.61\cosh d_{2}-1.44$
@@
$\log{\left(\kappa_{\rm L}\right)_{{\rm{SISSO}}}}=$ $1.53+0.0079\frac{{F\log\left({{C_{\rm{v}}}}\right)}}{{\sqrt[3]{{{C_{\rm{v}}}}}}}$
@@ II.3 Screening potential thermal insulators/conductors
Figure 7: $k$-means clustering and t-SNE 2D visualization. (a) Labeled training samples from AFLOW are clustered into four groups, each represented by a different color. (b) The same scatter plot as in (a), but the samples are colored according to their $\log(\kappa_{\rm L})$, with black pentagrams representing unlabeled candidates from the MP database. (c) and (d) are zoomed-in views showing the thermal insulator/thermal conductor candidates in the t-SNE space.
Table 3: The predicted results of $\kappa_{\rm L}$ based on KAN. The underlined materials have available DFT and MD calculation results.
@@
  • Excluding transition metal elements;
  • Number of atoms (nsites) $<$ 5;
  • Band gap $\in\left({0,1.5{\rm{eV}}}\right]$.
In Figure 7(a), all training samples are distinctly clustered into seven clusters, $C_{1}$ to $C_{4}$, with well-defined boundaries between clusters. Mapping the values of $\log(\kappa_{\rm L})$ onto the classification results in Figure 7(a), as shown in Figure 7(b), we observe that as t-SNE component 1 increases, $\log(\kappa_{\rm L})$ exhibits an upward trend. This allows us to identify $C_{1}$ as the cluster corresponding to thermal insulators and $C_{4}$ as the cluster representing thermal conductors. More importantly, the clustering results indicate that $\log(\kappa_{\rm L})$ can be accurately and qualitatively assessed using only the three features $F$, $K_{\rm s}$, and $\alpha$. Based on the labeled dataset, we obtained a CGCNN pre-trained model for the initial features and applied it to predict $F$, $K_{\rm s}$, and $\alpha$ for unlabeled samples from the MP database. These predicted features were then clustered in the same manner as the labeled samples. In Figure 7(b), we selected the 10 most promising thermal insulator/conductor candidates from clusters $C_{1}$ and $C_{4}$ (see Table 3). The zoomed-in views are shown in Figures 7(c) and (d), where these candidates are found in the same clusters as the labeled samples with the lowest/highest $\log(\kappa_{\rm L})$, respectively. Experimental results proved that only 4 out of the 10 candidates (CN2, CrB4, Be2C and KRbAu2) effectively utilize DFT and MD to calculate $C_{\rm v}$ and $\kappa_{\rm L}$, respectively. The subsequent validation process will also focus on these 4 feasible structures.
Figure 8: Validation of the KAN predictions. For (a) Be2C, (b) CN2 and (c) CrB4, a relaxation time of 5 ns is sufficient for the $\kappa_{\rm L}$ to converge. The deep red dashed line represents the $\kappa_{\rm L}$ predicted by the KAN model, while the light red band indicates the prediction range with a ±5% error margin. The $\kappa_{\rm L}$ calculated by MD is represented by the blue line, illustrating its convergence behavior with respect to relaxation time. (d) The blue circular scatter points and connecting lines represent the temperature dependence curve of $\kappa_{\rm L}$ for KRbAu2, while the red dashed line represents the $\kappa_{\rm L}$ predicted by the KAN model. (e) shows a histogram of the $\kappa_{\rm L}$ distribution in the training dataset.
Among the thermal conductor candidates, only Be2C belongs to the cubic crystal system. CN2 exhibits high thermal conductivity along the $z$-direction while being nearly insulating in the $x$- and $y$-directions. CrB4 demonstrates anisotropic thermal conductivity in all three directions. Following the approach in relevant studies [74, 75], we employed MD simulations to validate the $\kappa_{\rm L}$ values of these conductor candidates. For Be2C, calculations are required for only one direction due to its cubic symmetry. For CN2, we considered only $\kappa_{\rm L}$ along the $z$-axis. In the case of CrB4, the overall $\kappa_{\rm L}$ is determined as the average of the values computed along the three principal directions.
According to our MD calculations, the actual $\kappa_{\rm L}$ values of the candidate materials fall within distinct ranges ($\sim$30, $\sim$80, $\sim$150 W/m/K). As shown in Figure 8(e), samples with $\kappa_{\rm L}>100$ W/m/K account for only 0.47% of the entire training dataset. In such cases of data imbalance, model predictions for regions with scarce samples typically exhibit considerable uncertainty [76, 77].
KRbAu2 has the space group symbol $\textit{Fm}\overline{3}\textit{m}$, belonging to the cubic crystal system. For such isotropic thermal insulator candidates, it is generally more efficient and accurate to calculate $\kappa_{\rm L}$ using DFT [78, 79]. As shown in Figure 8(d), the DFT-calculated $\kappa_{\rm L}$ for this structure at 300K is 0.201 W/m/K, while the KAN-predicted value is 0.2169, correctly identifying it as a thermal insulator. In summary, the methods of $k$-means clustering and t-SNE visualization are also sufficient for qualitatively assessing the $\kappa_{\rm L}$ of materials, enabling high-throughput screening in combination with the pretrained model. Furthermore, the $\kappa_{\rm L}$ values predicted by KAN exhibit excellent agreement with both MD and DFT validation results. This demonstrates that the pretrained KAN model achieves not only reliable accuracy but also satisfactory extrapolation capability.
II.4 Tracing high/low LTC from the perspective of theoretical calculation
Among the validated candidates, considering the isotropy of Be2C and KRbAu2, we take these two as examples to explore the mechanisms influencing high and low $\kappa_{\rm L}$ based on DFT. Their crystal structures are shown in Figures 9(a) and (b), both belonging to the cubic system but exhibiting distinctly different phonon transport behaviors. Be2C adopts a [BeC4] tetrahedral configuration, where each C atom is coordinated with four Be atoms, with a bond length of approximately 1.64 ${\rm\AA}$. These [BeC4] tetrahedral units share edges, forming a continuous three-dimensional covalent network that imparts high rigidity and thermal stability to the crystal. In contrast, KRbAu2 consists of K, Rb, and Au, forming a metal-ion mixed-bond network. The Au atoms exist as [Au2] dimers, where each Au atom is connected via metallic bonds to form a framework structure. K and Rb act as cations filling the interstitial spaces, providing electrostatic stabilization. Due to the large atomic radii of K and Rb, the introduced steric hindrance effect increases lattice flexibility [80].
Figure 9: Based on DFT-based phonon thermal transport analysis. Crystal structures and the projected 2D electron localization function (ELF) diagram of (a) CrB4 and (b) KRbAu2. (c) is the -pCOHP & -IpCOHP for Be-C, K-Au and Rb-Au bonds. In (d) and (e), the left panels depict the phonon dispersion curves of CrB4 and KRbAu2, while the right panels present the corresponding atom-projected PDOS.
The effects of lattice rigidity and flexibility on $\kappa_{\rm L}$ primarily manifest in their influence on phonon group velocity $v_{g}$ and anharmonicity [81]. In Be2C, the strong covalent bonding between Be and C forms a highly rigid lattice, leading to smaller atomic displacements and lower lattice vibrational anharmonicity, thereby reducing three-phonon scattering (primarily suppressing Umklapp scattering) [82]. Additionally, such crystals typically exhibit high elastic moduli, resulting in steep phonon dispersion curves, which lead to higher phonon velocities and enhanced thermal conductivity [83]. In contrast, the softer bonds between K, Rb, and Au make the lattice more deformable. A more flexible lattice experiences larger thermal vibrations, inducing stronger three-phonon and four-phonon scattering, which significantly reduces $\kappa_{\rm L}$. Atomic mass also indirectly influences $\kappa_{\rm L}$ by affecting sound velocity. Be2C consists of light atoms, whereas KRbAu2 is composed entirely of heavy atoms. Specifically, heavy atoms tend to reduce group velocity, thereby exhibiting thermal insulation characteristics, whereas light atoms have the opposite effect [84, 85, 86].
Figure 10: Modal quantities of CrB4 and KRbAu2. (a-c) illustrate the frequency dependence of the average thermal conductivity $\bar{\kappa}_{\lambda}$, group velocity $|\mathbf{v}_{\lambda}|$, and phonon lifetime $\tau_{\lambda}$ for CrB4, respectively; (d-f) present the counterparts for KRbAu2. The scatters are colored according to the band index, with two distinct colormaps used to differentiate the two materials.
To further investigate the impact of chemical bonding on the $\kappa_{\rm L}$ of Be2C and KRbAu2, we employ the Negative Projected Crystal Orbital Hamilton Population (-pCOHP) to visualize their local bonding characteristics. In -pCOHP, positive values indicate bonding states, negative values represent antibonding states, and a value of zero signifies non-bonding interactions. As shown in Figure 9(c), the Fermi level $E_{\rm F}$ lies at 0 eV. For the Be-C bond, bonding states dominate below $E_{\rm F}$ down to -8.7 eV. Strong bonding implies greater lattice rigidity and a steeper potential energy surface [87], leading to higher optical branch vibration frequencies, which suppress phonon scattering and enhance thermal conductivity. This behavior is similar to classical high-$\kappa_{\rm L}$ materials such as diamond and SiC. In KRbAu2, the bonding characteristics differ between the two types of bonds: K-Au is bonding, while Rb-Au is antibonding. The antibonding nature of the Rb-Au bond softens the lattice, enhances phonon scattering, and reduces $\kappa_{\rm L}$. Figure 8(d) confirms the ultra-low $\kappa_{\rm L}$ of KRbAu2, suggesting that its phonon transport properties are primarily dictated by the Rb-Au bond. The -IpCOHP analysis further evaluates global bond strength. Despite K-Au being a bonding interaction, its bond strength remains weak, and the presence of an antibonding Rb-Au bond exacerbates phonon anharmonicity. Weak chemical bonds are often associated with strong phonon anharmonicity [88, 89], which is a key factor underlying the ultra-low $\kappa_{\rm L}$ of KRbAu2.
Figures 9(d) and (e) provide insights into the characteristics of low-energy optical branches and acoustic modes, supporting the previous hypotheses. In the phonon dispersion spectrum of Be2C, all phonon modes exhibit real frequencies, indicating dynamical stability at zero temperature with no tendency for structural collapse. The acoustic branches in the low-frequency region (0 $\sim$ 10 THz) show relatively high group velocities, suggesting potentially high phonon thermal conductivity. In contrast, the optical branches are primarily distributed in the high-frequency range (15 $\sim$ 30 THz) and exhibit a certain degree of flatness, indicating lower group velocities and minimal contribution to heat transport. The Projected Density of State (PDOS) analysis on the right side of Figure 9(d) further reveals that heat-carrying phonons are predominantly below 15 THz, mainly contributed by the heavier Be atoms, while the high-frequency optical branches involve both C and Be atoms. For KRbAu2, the phonon dispersion curve shows that all modes remain within the real frequency range, with no imaginary branches, confirming its dynamical stability at zero temperature. The acoustic branches are concentrated in the low-frequency region (0 $\sim$ 1 THz) and exhibit flat dispersions, resulting in low group velocities. The optical branches are distributed between 1 and 3.5 THz, with some modes displaying significant dispersion. According to the PDOS, the dominant vibrating atoms shift sequentially from Au to Rb to K as frequency increases, consistent with the mass distribution of these elements.
Figure 10(b) and (e) show the frequency dependence of the group velocity $\mathbf{v}_{g}$ at the center of the Brillouin zone for Be2C and KRbAu2, respectively. The group velocity of Be2C exceeds 10 km/s, while that of KRbAu2 does not exceed 1.1 km/s. This significant difference in group velocity can be attributed to differences in bonding and strength, which is consistent with the inference made from the steepness of their phonon spectra. Figure 10(c) and (f) describe the differences in phonon lifetimes between the two materials. In the low-frequency range below 10 THz, Be2C exhibits longer phonon lifetimes, with the longest exceeding 10 ps and the shortest being above 1 ps. In contrast, in KRbAu2, the phonon lifetime decreases sharply with increasing frequency, with $\tau_{\lambda}$ approaching 0.1 ps at $\upsilon_{\lambda}=2{\rm THz}$. It is noteworthy that for certain specific low-frequency modes, the phonon lifetime $\tau_{\lambda}$ of KRbAu2 exceeds that of Be2C, which may be due to special scattering suppression effects, e.g., weaker three-phonon scattering [90]. However, overall, Be2C has a longer and more uniform phonon lifetime, which is favorable for long-range heat transport, while the shorter phonon lifetime of KRbAu2 suppresses $\kappa_{\rm L}$.